if (a-b)=7 and ab=9. then a^2+b^2=_____
1
Expert's answer
2014-06-23T02:19:59-0400
To find the answer using the next formula: (a-b)2=a2-2ab+b2.
So we know that (a-b)=7, from this we have (a-b)2=49.
On the other hand 49=a2-2ab+b2. Substitute here ab=9. We obtain: 49=a2-2*9+b2.

Thus, a2+b2=49+18=67.

Answer. 67.